# Asymptotically good CSS-T codes exist

### v1

@@ A CSS code $(C_{1},C_{2})_{\mathrm{CSS}}$ is called a CSS-T code if the followings hold:
A reformulation of the above conditions in terms of the Schur product of the ingredient classical codes was proved in [9] and we recalled below.
Theorem 2 ([9, Theorem 2.3]).
A pair of binary linear codes $C_{1}$ and $C_{2}$ generates a CSS-T code if and only if $C_{1}^{\perp}+C_{1}^{\star 2}\subseteq C_{2}^{\perp}$.
It is worth noting that if we consider a pair of linear codes $C_{2}\subseteq C_{1}$ giving a CSS code, we obtain for free the inclusion $C_{1}^{\perp}\subseteq C_{2}^{\perp}$. Therefore, each given CSS code $(C_{1},C_{2})_{\mathrm{CSS}}$, is a CSS-T code if and only if $C_{1}^{\star 2}\subseteq C_{2}^{\perp}$.
Some known constructions and several examples of CSS-T codes are provided in [1, 8, 9, 10, 32, 33]. All currently known families of CSS-T codes have either a vanishing rate or a vanishing relative distance. An interesting open problem, which will be answered in the next section, is to construct an infinite family of asymptotically good CSS-T codes with non-vanishing rate and relative distance [32].
2 Constructing CSS-T codes from CSS codes
In this section, we present a method to construct a CSS-T code from any given CSS code. This allows us to prove in Theorem 4 and Corollary 1 the existence of families of asymptotically good CSS-T codes. In particular, our construction preserves the LDPC property of the original CSS code, enabling us to show in Theorem 5 and Corollary 2 the existence of asymptotically good quantum LDPC codes that are CSS-T.
Our approach is based on a technique that doubles the ingredient pair of classical codes. This entire section builds up to the following result.
Theorem 3.
In the remainder of this section, we provide more details about our construction and prove the above mentioned result.
@@ 2 Constructing CSS-T codes from CSS codes
By definition, $C_{2}^{N_{\phi}}\subseteq C_{1}^{N_{\phi}}$. Hence, $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)$ is a CSS code if and only if $C_{1}^{N_{\phi}}$ and $C_{2}^{N_{\phi}}$ are linear codes, that is, if and only if for any $x,y\in C_{1}$ we have $(x,\phi(x))+(y,\phi(y))=(x+y,\phi(x)+\phi(y))=(x+y,\phi(x+y))$, where the second equality follows by the definition of $C_{1}^{N_{\phi}}$. Thus, $\phi$ must be linear on $C_{1}$. Now, $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)_{\mathrm{CSS}}$ is CSS-T if and only if $\left({C_{1}^{N_{\phi}}}\right)^{\star 2}\subseteq\left({C_{2}^{N_{\phi}}}\right)^{\perp}$ by Theorem 2. So $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)_{\mathrm{CSS}}$ is CSS-T if and only if for each $x,y\in C_{1}$ and each $z\in C_{2}$ we have
@@
$=\sum_{i=1}^{n}x_{i}y_{i}z_{i}+\sum_{i=1}^{n}\phi(x)_{i}\phi(y)_{i}\phi(z)_{i}=0.$
The last equality holds if and only if ${\textrm{w}t}(x\star y\star z)+{\textrm{w}t}(\phi(x)\star\phi(y)\star\phi(z))\equiv 0\pmod{2}$. This completes the proof. ∎
@@
Clearly, the two CSS codes have the same length and the same dimension. So, we only need to deal with the minimum distance. Let $\Phi=(I,\phi)$ be the permutation of $\mathbb{F}_{2}^{2n}$ that leaves the first $n$ coordinates unchanged and applies $\phi$ to the second $n$ coordinates. Then $\Phi\left(C_{1}^{N_{I}}\right)=C_{1}^{N_{\phi}}$ and $\Phi\left(C_{2}^{N_{I}}\right)=C_{2}^{N_{\phi}}$. Therefore, $\Phi\left(C_{1}^{N_{I}}\setminus C_{2}^{N_{I}}\right)=C_{1}^{N_{\phi}}\setminus C_{2}^{N_{\phi}}$. Hence, both CSS codes have the same $X$-distance.
Since $\Phi$ is a permutation, we have $\Phi\left(\left({C_{i}^{N_{I}}}\right)^{\bot}\right)=\left(\Phi\left(C_{i}^{N_{I}}\right)\right)^{\bot}$ for $i\in\{1,2\}$. Thus $\Phi\left(\left({C_{i}^{N_{I}}}\right)^{\bot}\right)=\left(\Phi\left(C_{i}^{N_{I}}\right)\right)^{\bot}=\left(C_{i}^{N_{\phi}}\right)^{\bot}$ and $\Phi\left({\left(C_{2}^{N_{I}}\right)}^{\bot}\setminus{\left(C_{1}^{N_{I}}\right)}^{\bot}\right)={\left(C_{2}^{N_{\phi}}\right)}^{\bot}\setminus{\left(C_{1}^{N_{\phi}}\right)}^{\bot}$ for $i\in\{1,2\}$. Hence, both CSS codes have the same $Z$-distance. ∎
@@
which we will denote simply by $C^{N}$.
@@
Other choices of $\phi$ are possible and may lead to CSS-T codes with different parameters. As an example, consider the map $\phi:x\mapsto x+a$ for some $a\in C_{1}\cap(C_{1}^{\star 2})^{\perp}$. Finding maps $\phi$ giving CSS-T codes with better parameters than the choice $\phi=I$ is an interesting open question. We return to this discussion in the conclusion.
By Proposition 2 we know that $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS-T}}$ has length $2n$ and dimension $k=\dim(C_{1})-\dim(C_{2})$. Hence, in what follows, we focus on studying the minimum distance of this code. To this end, we set $d^{N}=\min\left\{d_{X}^{N},d_{Z}^{N}\right\}$ for the CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS-T}}$, where $d_{X}^{N}=\min\left\{{\textrm{w}t}\left(C_{1}^{N}\setminus C_{2}^{N}\right)\right\}$ and $d_{Z}^{N}=\min\left\{{\textrm{w}t}\left((C_{2}^{N})^{\perp}\setminus(C_{1}^{N})^{\perp}\right)\right\}$.
Lemma 1.
The CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS}}$ satisfies $d_{X}^{N}=2d_{X}$.
@@
Let $x^{N}=(x,x)$ be the minimum weight codeword of $C_{1}^{N}\setminus C_{2}^{N}$, where $x\in C_{1}$. Note that $x\in C_{1}\setminus C_{2}$ as otherwise $x^{N}\in C_{2}^{N}$. The fact that $x^{N}$ has the minimum weight implies that $x$ has to have weight $d_{X}$. Therefore, ${\textrm{w}t}(x^{N})=2d_{X}$ and we have $d_{X}^{N}=2d_{X}$. ∎
To calculate $d_{Z}^{N}$, we need the following technical lemma.
Lemma 2.
The followings hold.
  1. 1.
    Let $x=(x_{1},x_{2})\in\left(C_{2}^{N}\right)^{\perp},$ for some $x_{i}\in\mathbb{F}_{2}^{n}$. Then $x^{\prime}=(x_{1}+x_{2},0)\in\left(C_{2}^{N}\right)^{\perp}$. In particular, $x_{1}+x_{2}\in C_{2}^{\bot}$ and ${\textrm{w}t}(x^{\prime})\leq{\textrm{w}t}(x)$.
  2. 2.
    For each $x=(a,a)\in\left(C_{2}^{N}\right)^{\perp}$, we have $x\in\left(C_{1}^{N}\right)^{\perp}$.
  3. 3.
    Let $x_{1},x_{2}\in\mathbb{F}_{2}^{n}$. Then $x_{1}+x_{2}\in C_{1}^{\perp}$ if and only if $x^{\prime}=(x_{1}+x_{2},0)\in\left(C_{1}^{N}\right)^{\perp}$.
Proof.
(1) For each $y=(a,a)\in C_{2}^{N}$ we have
$x\cdot y=x_{1}\cdot a+x_{2}\cdot a=(x_{1}+x_{2})\cdot a=x^{\prime}\cdot y=0.$
The fact that $x_{1}+x_{2}\in C_{2}^{\bot}$ and ${\textrm{w}t}(x^{\prime})\leq{\textrm{w}t}(x)$ follows immediately.
(2) Let $y=(b,b)\in C_{1}^{N}$. Then $x\cdot y=a\cdot b+a\cdot b=0$. Hence $x\in\left(C_{1}^{N}\right)^{\perp}$.
(3) The statement is proved by the following sequence of equivalences:
$x_{1}+x_{2}\in C_{1}^{\perp}$ $\iff\forall a\in C_{1},(x_{1}+x_{2})\cdot a=0\iff\forall y=(a,a)\in C_{1}^{N},x^{\prime}\cdot y=0$
$\iff x^{\prime}=(x_{1}+x_{2},0)\in\left(C_{1}^{N}\right)^{\perp}.$
Next, we use the above lemma to compute $d_{Z}^{N}$.
Proposition 4.
The CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS}}$ satisfies $d_{Z}^{N}=d_{Z}$.
Proof.
Let $x=(x_{1},x_{2})\in\left(C_{2}^{N}\right)^{\perp}\setminus\left(C_{1}^{N}\right)^{\perp}$ for some $x_{1},x_{2}\in\mathbb{F}_{2}^{n}$ be a minimum weight vector. Since $d_{Z}^{N}={\textrm{w}t}\left(\left(C_{2}^{N}\right)^{\perp}\setminus\left(C_{1}^{N}\right)^{\perp}\right)$, Lemma 2(2) implies that $x_{1}\neq x_{2}$. Next, Lemma 2(2) implies that $x^{\prime}=(x_{1}+x_{2},0)\in\left(C_{2}^{N}\right)^{\perp}$. Moreover, the fact that $x\notin\left(C_{1}^{N}\right)^{\perp}$ implies the existence of $c\neq 0\in C_{1}$ such that $x\cdot(c,c)=x_{1}\cdot c+x_{2}\cdot c\neq 0$. Therefore, $x^{\prime}\cdot(c,c)\neq 0$ which implies $x^{\prime}\notin\left(C_{1}^{N}\right)^{\perp}$. Now, Lemma 2(3) implies that $x_{1}+x_{2}\in\left(C_{2}^{\perp}\setminus C_{1}^{\perp}\right)$. This shows that
$d_{Z}^{N}={\textrm{w}t}(x)\geq{\textrm{w}t}(x^{\prime})={\textrm{w}t}(x_{1}+x_{2})\geq d_{Z}.$ (3)
For the other inequality, let $x$ be a vector of minimum weight in $C_{2}^{\perp}\setminus C_{1}^{\perp}$. Then, observe that $(x,0)\in\left(C_{2}^{N}\right)^{\perp}\setminus\left(C_{1}^{N}\right)^{\perp}$ as for each $(y,y)\in C_{2}^{N}$, where $y\in C_{2}$, we have $(x,0)\cdot(y,y)=x\cdot y=0.$ The fact that $(x,0)\not\in(C_{1}^{N})^{\perp}$ follows from Lemma 2(3). Thus
${\textrm{w}t}((x,0))=d_{Z}\geq d_{Z}^{N}.$ (4)
Finally, Equations (3) and (4) complete the proof. ∎
Finally, we are ready to prove Theorem 3.
Proof of Theorem 3.
Let $(C_{1},C_{2})_{\mathrm{CSS}}$ have parameters $[\![n,k,d]\!]$. The combination of Propositions 2 and 4 and Lemma 1 shows that the CSS-T family $(C_{1}^{N},C_{2}^{N})$ has parameters $[\![2n,k,\geq d]\!]$. This concludes the proof. ∎
Remark 2.
Let $(C_{1},C_{2})_{\mathrm{CSS}}$ be a CSS code with the $X$ and $Z$ distances $d_{X}$ and $d_{Z}$, respectively.
  1. 1.
    When $d_{X}<d_{Z}$ the CSS-T code of Theorem 3 has a strictly larger minimum distance than the initial CSS code. On the other hand, if $d_{X}>d_{Z}$ the CSS-T code has the same minimum distance as the initial CSS code.
  2. 2.
    When $d_{X}\neq d_{Z}$, one can apply the result of Theorem 3 to construct CSS-T code with double length, improved minimum distance and the same dimension of the ingredient CSS code. In particular, if $d_{X}<d_{Z}$ in a CSS code $(C_{1},C_{2})_{CSS}$, then applying Theorem 3 directly to this code gives the desired CSS-T code. In the case when $d_{X}>d_{Z}$, then one can consider $C_{1}^{\prime}=C_{2}^{\perp}$ and $C_{2}^{\prime}=C_{1}^{\perp}$, after exchanging the $X$ and $Z$ stabilizers in the CSS code. Then, applying Theorem 3 to $C_{2}^{\prime}\subseteq C_{1}^{\prime}$ gives the desired result.
  3. 3.
    If $d_{X},d_{Z}>2$, then the CSS-T code of Theorem 3 is a degenerate quantum code. In particular, it contains all weight two codewords in the form $(a,a)$, where $a\in\mathbb{F}_{2}^{n}$ is a weight one vector.
Table 1 provides a list of examples of binary CSS-T codes that we found after applying the construction of Theorem 3 to binary linear cyclic codes.
@@
Table 1: Small length binary CSS-T codes from cyclic codes. All the CSS-T codes are $Z$-degenerate. The codes shown by $\ast$ are also $X$-degenerate.
2.1 Asymptotically good families of CSS-T codes
Recall that the rate and relative distance of a binary quantum code with parameters $[\![n,k,d]\!]$ are defined as $R=\frac{k}{n}$ and $\delta=\frac{d}{n}$, respectively. In the literature there exist several infinite families of quantum CSS codes with non-vanishing asymptotic rate and relative distance, commonly known as asymptotically good quantum codes, see for example [2, Theorem 1], [11, Theorem 1.2], and [28, Theorem 2]. The existence of such families combined with our Theorem 3 allow us to construct asymptotically good CSS-T codes.

### v2

@@ A CSS code $(C_{1},C_{2})_{\mathrm{CSS}}$ is called a CSS-T code if the following hold:
Importantly, a CSS-T code is defined as a CSS code for which a transversal application of $T$ gates at the physical level, i.e., $T^{\otimes n}$, preserves the codespace. In this sense, the resulting logical operator for CSS-T codes can be either the logical identity or the logical $T$ after the application of $T^{\otimes n}$ at the physical level. A reformulation of the above conditions in terms of the Schur product of the ingredient classical codes was proved in [14], and we recall below.
Theorem 2 ([14, Theorem 2.3]).
A pair of binary linear codes $C_{1},C_{2}$ generates a CSS-T code if and only if $C_{1}^{\perp}+C_{1}^{\star 2}\subseteq C_{2}^{\perp}$.
If we consider a pair of linear codes $C_{2}\subseteq C_{1}$ giving a CSS code, we obtain the inclusion $C_{1}^{\perp}\subseteq C_{2}^{\perp}$. Therefore, each given CSS code $(C_{1},C_{2})_{\mathrm{CSS}}$, is a CSS-T code if and only if $C_{1}^{\star 2}\subseteq C_{2}^{\perp}$. Note also that for each $x,y\in C_{1}$, the support of the Schur product $x\star y$ is the set of coordinate positions where both $x$ and $y$ are nonzero (also known as the intersection of corresponding stabilizers). Note that a CSS-T code may have stabilizer generators with negative signs. As mentioned in [16, Remark 3], applying a Pauli correction can transform the code space into the form of Equation Eq. 1. Thus, without loss of generality, we assume this encoding.
Some constructions and examples of CSS-T codes are provided in [12, 14, 15, 16, 17]. All currently known families of CSS-T codes have either a vanishing rate or a vanishing relative distance. An interesting open problem, which will be solved in the next section, is constructing an infinite family of asymptotically good CSS-T codes with non-vanishing rate and relative distance [16].
II-D Triorthogonal codes
Quantum triorthogonal codes are a subclass of binary CSS-T codes which can realize the logical $T$ gate in a transversal manner (up to Pauli corrections). Many magic state distillation protocols rely on such codes, see e.g., [11, 24, 25]. Here, we briefly recall the definition of binary triorthogonal codes.
Definition 2 (Triorthogonal Matrices, [11]).
A binary matrix $G$ of size $m\times n$ is called triorthogonal if and only if the supports of any pair and any triple of its distinct rows have even overlap, i.e.,
$\sum_{j=1}^{n}G_{a,j}G_{b_{j}}\equiv 0\mod{2}$
for all pairs of rows $1\leq a<b\leq m$ and
$\sum_{j=1}^{n}G_{a,j}G_{b_{j}}G_{c_{j}}\equiv 0\mod{2}$
for all triples of rows $1\leq a<b<c\leq m.$
We call the linear code generated by a triorthogonal matrix a triorthogonal subspace. We can write a triorthogonal matrix as $G=[G_{1},G_{0}]^{T}$, where $G_{1}$ consists of odd-weight rows, and $G_{0}$ consists of even-weight rows.
Theorem 3 (Triorthogonal Codes [11]).
Let $G=[G_{1},G_{0}]^{T}$ be a triorthogonal matrix, where $G_{1}$ consists of all $k$ odd-weight rows. Let $[\![n,k]\!]$ be the CSS code of $(C_{1},C_{2})$ formed when $G_{0}$ and $G$ (with $G\neq G_{0}$) are taken as the generator matrices of $C_{2}$ and $C_{1}$, respectively. Then applying transversal $T$ to the physical qubits realizes the transversal logical $T$ (up to Pauli corrections).
III Asymptotically good CSS-T codes exist
The main goal of this section is to show the existence of asymptotically good CSS-T codes, thus answering the open question asked in [16]. We start by presenting a method to construct a CSS-T code from any given CSS code using an auxiliary map $\phi$. When $\phi$ is the identity map, we recover the construction and parameters presented in [1]. Moreover, when $\phi$ is a permutation map, we show that this construction yields the same parameters as in the identity case. The CSS-T codes constructed using this method allow us to prove in Theorem 5 and Corollary 1 the existence of families of asymptotically good CSS-T codes. In particular, this construction preserves the LDPC property of the original CSS code, enabling us to show in Theorem 6 and Corollary 2 the existence of asymptotically good quantum LDPC codes that are CSS-T.
@@ III Asymptotically good CSS-T codes exist
By definition, $C_{2}^{N_{\phi}}\subseteq C_{1}^{N_{\phi}}$. Hence, $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)$ is a CSS code if and only if $C_{1}^{N_{\phi}}$ and $C_{2}^{N_{\phi}}$ are linear codes, that is, if and only if for any $x,y\in C_{1}$ we have $(x,\phi(x))+(y,\phi(y))=(x+y,\phi(x)+\phi(y))=(x+y,\phi(x+y))$, where the second equality follows by the definition of $C_{1}^{N_{\phi}}$. Thus, $\phi$ must be linear on $C_{1}$. Now, $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)_{\mathrm{CSS}}$ is CSS-T if and only if $\left({C_{1}^{N_{\phi}}}\right)^{\star 2}\subseteq\left({C_{2}^{N_{\phi}}}\right)^{\perp}$ by Theorem 2. So $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)_{\mathrm{CSS}}$ is CSS-T if and only if for each $x,y\in C_{1}$, $z\in C_{2}$ we have
@@
$=\sum_{i=1}^{n}x_{i}y_{i}z_{i}+\sum_{i=1}^{n}\phi(x)_{i}\phi(y)_{i}\phi(z)_{i}$
$=0.$
The last equality holds if and only if ${\textrm{w}t}(x\star y\star z)+{\textrm{w}t}(\phi(x)\star\phi(y)\star\phi(z))\equiv 0\pmod{2}$. ∎
@@
Clearly, the two CSS codes have the same length and dimension. So, we only need to deal with the minimum distance. Let $\Phi=(I,\phi)$ be the permutation of $\mathbb{F}_{2}^{2n}$ that leaves the first $n$ coordinates unchanged and applies $\phi$ to the second $n$ coordinates. Then $\Phi\left(C_{1}^{N_{I}}\right)=C_{1}^{N_{\phi}}$ and $\Phi\left(C_{2}^{N_{I}}\right)=C_{2}^{N_{\phi}}$. Therefore, $\Phi\left(C_{1}^{N_{I}}\setminus C_{2}^{N_{I}}\right)=C_{1}^{N_{\phi}}\setminus C_{2}^{N_{\phi}}$. Hence, both CSS codes have the same $X$-distance. Since $\Phi$ is a permutation, we have $\Phi\left(\left({C_{i}^{N_{I}}}\right)^{\bot}\right)=\left(\Phi\left(C_{i}^{N_{I}}\right)\right)^{\bot}$ for $i\in\{1,2\}$. Thus $\Phi\left(\left({C_{i}^{N_{I}}}\right)^{\bot}\right)=\left(\Phi\left(C_{i}^{N_{I}}\right)\right)^{\bot}=\left(C_{i}^{N_{\phi}}\right)^{\bot}$ and $\Phi\left({\left(C_{2}^{N_{I}}\right)}^{\bot}\setminus{\left(C_{1}^{N_{I}}\right)}^{\bot}\right)={\left(C_{2}^{N_{\phi}}\right)}^{\bot}\setminus{\left(C_{1}^{N_{\phi}}\right)}^{\bot}$ for $i\in\{1,2\}$. Hence, both CSS codes have the same $Z$-distance. ∎
@@
which we will denote simply by $C^{N}$. This code coincides with the construction proposed in [1, Section VII], when taking $M=2$. Hence, relying on this previous work, we get the following result.
Theorem 4.
Proof.
Consider the CSS-T code $(C_{1}^{N},C_{2}^{N})$. By Proposition 2 we know that $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS-T}}$ has length $2n$ and dimension $k$. The bound on the minimum distance follows from [1, Theorem 13]. ∎
@@
Other choices of $\phi$ are possible and may lead to CSS-T codes with different parameters. As an example, consider the map $\phi$ defined by $\phi(x)=x+a$ for some $a\in C_{1}\cap(C_{1}^{\star 2})^{\perp}$ and $x\in\mathcal{B}$ for some basis $\mathcal{B}$ of $C_{1}$. Finding maps $\phi$ giving CSS-T codes with better parameters than the choice $\phi=I$ or a permutation is an interesting open question. We will return to it in the conclusion.
The next proposition shows that applying transversal $T$ to all the physical qubits of CSS-T codes satisfying Theorem 4 implies a logical identity operator.
Proposition 4.
Let $(C_{1},C_{2})_{\mathrm{CSS}}$ be a CSS-T pair. If $C_{1}^{\star 2}\subseteq C_{1}^{\perp}$, then applying transversal $T$ to all the physical qubits implies a logical identity operator. In particular, all the CSS-T codes of Theorem 4 satisfy this property.
@@
The condition $C_{1}^{\star 2}\subseteq C_{1}^{\perp}$ implies that $C_{1}\subseteq C_{1}^{\perp}$, $C_{1}\star C_{2}\subseteq C_{1}^{\perp}$, $C_{1}\star C_{1}\subseteq C_{1}^{\perp}$, and $C_{2}\star C_{2}\subseteq C_{1}^{\perp}$, which are conditions 1–4 of [17, Theorem 12], from which the result follows. For the second part, let $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS}}$ be a CSS-T code of Theorem 4, where $C_{i}^{N}=(C_{i},C_{i})$, $i\in\{1,2\}$, for two length $n$ binary codes $C_{1}$ and $C_{2}$. For each $x,y\in C_{1}$ and each $z\in C_{1}$ we have
$((x,x)\star(y,y))\cdot(z,z)$ $=(x\star y)\cdot z+(x\star y)\cdot z=0.$
Thus $(C_{1}^{N})^{\star 2}\subseteq(C_{1}^{N})^{\perp}$ and this completes the proof. ∎
Table I provides a list of examples of binary CSS-T codes that we found after applying the construction of Theorem 4 to binary linear cyclic codes.
@@
TABLE I: Small length binary CSS-T codes from cyclic codes. All the CSS-T codes are $Z$-degenerate. The codes shown by $\ast$ are also $X$-degenerate.
III-A Asymptotically good families of CSS-T codes
Recall that the rate and relative distance of a binary quantum code with parameters $[\![n,k,d]\!]$ are defined as $R=\frac{k}{n}$ and $\delta=\frac{d}{n}$, respectively. In the literature, there exist several infinite families of quantum CSS codes with non-vanishing asymptotic rate and relative distance, commonly known as asymptotically good quantum codes, see for example [26, 27, 28]. The existence of such families, combined with our Theorem 4, allows us to construct asymptotically good CSS-T codes.